# Outcome-Refining Process Supervision for Code Generation

### v1

@@ 3.3 Execution-Guided Process Reward Model
4 Experiments
LBPP HumanEval MBPP
(Matton et al., 2024) (Chen et al., 2021b) (Austin et al., 2021)
# Test Problems 162 164 257
# Unit Tests 5.1 6.5 3.0
Solution Length§ 627 / 3039 169 / 622 130 / 589
Contamination New Dataset 18.9% 20.8%
Task Type Algorithms Func. Completion Basic Prog.
4.1 Experimental Setup
Datasets. We evaluate on 3 programming benchmarks as shown in Table 2. LBPP is a recent complex programming dataset manually curated by human experts with competitive programming experience. HumanEval and MBPP are popular code generation benchmarks but could be too easy for current LLMs (Matton et al., 2024). Moreover, a significant proportion of their data is leaked in popular pre-training corpora (Riddell et al., 2024). To ensure reproducibility, we report our detailed hyperparameters in  Appendix A, we also open-source all our code and scripts in the aforementioned URL.
Baselines. For outcome supervision, Reflexion (Shinn et al., 2024) is a recent self-improvement approach that leverages execution feedback for code refinement. LDB (Zhong et al., 2024) enhances this by incorporating execution outcomes, debugger outputs, and intermediate variable values to iteratively fix solutions. For test-time scaling, we implement Best-of-N sampling, which generates multiple solutions and selects the best based on test outcomes. Due to lack of existing process supervision methods for code generation, we implement a similar approach from mathematical reasoning (Luo et al., 2024) ourselves for comparison and show in ablation studies. As some methods require access to unit tests from benchmarks, we mark them as (w/ T) in our tables.
Evaluation Metrics. We primarily report Pass@1, while also tracking the average percentage of passed test cases and successful compilations to evaluate program correctness in detail. Additionally, we analyze code quality through complexity measures and resource utilization metrics including execution time, memory usage, and cyclomatic complexity, providing a comprehensive view beyond mere correctness. Detailed descriptions of these metrics are provided in Table 5.
@@ 4.2 Main Results
Particularly noteworthy is the performance boost when models have access to actual unit tests from test datasets (without access to solutions). All models show drastic improvements on all metrics in this setting. For instance, Qwen-7B achieves 77.8% Pass@1 on LBPP and 96.3% on HumanEval with test case access, compared to 59.9% and 79.9% without. This suggests that while our self-generated test cases may be relatively weak, given feedback for higher quality test cases, models can effectively guide themselves through the reasoning process to generate significantly better code.
@@
Table 3: Ablation Study Results. - Execution: Remove execution feedback from our framework. - Reasoning: Remove in-depth reasoning process. Every metric is in percentage.
@@
4.3 Component-wise Ablation Study
We conducted experiments on the challenging LBPP dataset using the Qwen-7B model to investigate the importance of incorporating execution outcomes and reasoning during the exploration process. The results are presented in  Table 3.
When the model is unable to access execution outcomes during exploration, Pass@1 decreases by 16.1%. This highlights the critical role of real-environment execution in guiding the model to generate correct solutions. It is relatively challenging for LLMs to predict the execution outcomes of a given piece of code (Jain et al., 2024). Integrating these results into the exploration process ensures that the model benefits from concrete and actionable feedback.
Conversely, omitting extensive reasoning during exploration results in a 4.3% decrease in Pass@1. Although this reduction is smaller compared to the absence of execution results, reasoning remains a vital component for performance enhancement. Reasoning enables the model to iteratively refine its approach based on feedback, addressing issues that may not be resolved through execution feedback alone. This iterative refinement is crucial, particularly for solving complex problems where execution feedback alone may be insufficient.
@@ 4.4 Analysis of Process Reward Model
Our framework uses an implicit process reward model (PRM), which provides outcome-level supervision during beam search without additional training. Most existing work on process supervision generates line-level process reward signals and relies on explicitly trained PRMs. This motivates us to explore two questions: (1) Are outcome-level rewards more effective than line-level rewards? (2) Is an implicit PRM that does not require training better than an explicitly trained PRM?
To address these questions, we conduct experiments on the LBPP dataset using the Qwen-7B model. For the outcome-level process supervision method, the implementation details are consistent with the corresponding parts of ORPS. For the line-level method, the model generates step-by-step thoughts for the coding problem, with numerical process rewards assigned to each step. The final code is then generated based on the best thought trace. For methods requiring explicit training, we randomly select half of the LBPP dataset as a training set to avoid data leakage. To simulate human-annotated process feedback, we filter data from GPT-4’s outputs.
Results in Table 4 confirm that our framework substantially outperforms the other three method, validating our design choices. Overall, outcome-level reward signals prove to be more effective than line-level signals. Intuitively, line-level signals can only provide feedback for incomplete thought processes, which undoubtedly lack more information compared to the outcome-level.
Additionally, the implicit PRM shows greater effectiveness than the explicit PRM. This suggests that external process supervision feedback may not always be reliable. We consider that LLMs already have strong self-reflection capabilities and only require execution outcomes to activate this ability. This also indicates that spending extra data and time on training reward models might be unnecessary.
@@ 4.5 Scaling Analysis
In addition, we wanted to explore how the performance of ORPS changes as reasoning overhead increases. For comparison, we chose BoN as the baseline. This is because BoN allows easy control of reasoning overhead with linear growth. We conducted experiments on two models using the most challenging LBPP dataset. The results are shown in Figure 4. With the same model, ORPS improved much faster as reasoning overhead increased. This shows that ORPS has strong scaling potential. It can effectively use more computational resources to improve reasoning. In comparison, BoN showed slower improvements, suggesting it does not fully utilize the increased reasoning capacity.
@@ 4.6 Case Studies
Figure 4: Performance vs. Inference Budget. Pass@1 scores on LBPP with varying inference budgets. Our method maintains superior performance across different computational constraints.
Through detailed case studies, we demonstrate how our framework enhances code generation by improving reasoning. As shown in Appendix C, the response generated by the traditional CoT method for the Minimum Greatest Common Divisor problem in LBPP demonstrates that while the model provides a detailed step-by-step thought process during solution generation, the complexity of the task results in an imperfect code implementation. For instance, in CoT’s approach, the reliance on nested loops and pairwise GCD calculations introduces inefficiencies and fails to address scalability for larger datasets. Similarly, our method’s initial implementation demonstrates a lack of robustness in handling edge cases and unnecessary redundancies in subset formation.
@@
Figure 5: Performance by Problem Class. Top-20 problem classes in LBPP showing success rates and unsolved cases for our method vs baseline.
@@ 5 Conclusion
In this study, we introduced Outcome-Refining Process Supervision (ORPS), a novel paradigm for enhancing code generation through structured reasoning and execution-driven feedback. By leveraging a tree-structured exploration framework, ORPS facilitates diverse solution trajectories, enabling models to refine both theoretical reasoning and practical implementation simultaneously. Our approach demonstrated significant improvements in correctness and efficiency across various benchmarks, including an average Pass@1 increase of 26.9% and a 42.2% reduction in runtime, outperforming traditional outcome-based and process-based supervision methods.
Key findings reveal that structured reasoning space and concrete feedback signals are pivotal for solving complex programming tasks. ORPS proved effective even with smaller models, underscoring the importance of reasoning capabilities over mere model scaling. Furthermore, our framework’s reliance on execution feedback eliminates the need for expensive, annotated training data, making it a cost-efficient alternative.
These contributions highlight the potential of process supervision to enhance complex problem-solving abilities in LLMs. Future work could extend this framework to other domains requiring rigorous reasoning and verification. By bridging the gap between reasoning quality and execution fidelity, ORPS paves the way for more autonomous and adaptive systems in computational intelligence.
6 Limitations
While ORPS  demonstrates significant advancements in code generation, it has notable limitations. First, its reliance on execution outcomes as primary feedback restricts its applicability to tasks where execution is feasible and well-defined. Ambiguous problem descriptions or creative tasks beyond executable code can hinder the framework’s performance. Additionally, this reliance may lead to a bias toward solutions optimizing immediate execution success, potentially overlooking broader algorithmic considerations.
The tree-structured beam search also introduces significant computational overhead, requiring substantial memory and processing power. While scalable with increased resources, this approach may be impractical for real-time or resource-constrained applications, limiting its utility in some environments.
Finally, the quality of ORPS depends heavily on the underlying language model. If the model lacks domain knowledge or reasoning ability, it may generate incoherent reasoning chains or fail to explore promising solutions effectively, especially when feedback mechanisms rely solely on execution outcomes without additional validation.
7 Ethical Considerations
While ORPS advances code generation capabilities, we acknowledge important ethical considerations. Like all LLM-based systems, our framework inherits potential biases from the underlying models. However, our execution-guided approach provides an advantage: concrete verification helps detect and mitigate certain failure modes through objective testing. The framework’s ability to explore multiple solution paths also reduces the risk of being stuck with problematic implementations.
We emphasize that ORPS is designed as a development aid rather than a replacement for human programmers. The execution feedback and reasoning traces make the code generation process more transparent and auditable. We encourage using this framework in conjunction with established software development practices, including code review and testing, particularly for applications in sensitive domains.
@@ A.4 Model Training Configuration
  • Training Framework: llamafactory (Zheng et al., 2024)
@@ A.5 Hardware Environment
Appendix B AI Usage in Code Development
@@
Appendix C Example Model Outputs

### v2

@@ 4.2 Main Results
Particularly noteworthy is the performance boost when models have access to gold unit tests from test datasets (without access to solutions). All models show drastic improvements on all metrics in this setting. For instance, Qwen-7B achieves 77.8% Pass@1 on LBPP and 96.3% on HumanEval with test case access, compared to 59.9% and 79.9% without. This suggests that while our self-generated test cases may be relatively weak, given feedback for higher quality test cases, models can effectively guide themselves through the reasoning process to generate significantly better code.
@@
Table 3: Ablation Study Results. - Execution: Remove execution from our framework. - Reasoning: Remove reasoning process. Every metric is in percentage.
@@
4.3 Ablation Study
We conducted experiments on the challenging LBPP dataset using the Qwen-7B model to investigate the importance of two key components in the exploration process:
  1. 1.
    Execution Outcomes. Studies whether the model can access execution results to guide solution refinement. Without execution feedback, the model must rely solely on internal reasoning to assess correctness.
  2. 2.
    Reasoning. Investigates whether the model should explicitly perform reasoning before generating new code at each step. Without reasoning, the model directly generates new code based only on past code and execution results.
The results are presented in  Table 3. When the model is unable to access execution outcomes during searching, Pass@1 decreases by 16.1%. This highlights the critical role of environment feedback in guiding the model to generate correct solutions. Since LLMs struggle to accurately predict execution outcomes for a given piece of code (Jain et al., 2024), incorporating execution results ensures that the model benefits from concrete feedback.
Omitting reasoning during searching results in a 4.3% decrease in Pass@1. Reasoning enables the model to iteratively refine its approach based on feedback, addressing issues that may not be resolved through execution feedback.
@@ 4.4 Analysis of Process Reward Model
Our framework challenges the necessity of training Process Reward Models (PRMs) by combining outcome rewards with reasoning to work as process rewards. While most prior work on process supervision (Luo et al., 2024) assumes that specially trained PRMs are required to guide reasoning during inference, we demonstrate that execution feedback as verifiable rewards, combined with existing LLMs’ reasoning capabilities, can generate high-quality process rewards. This motivates us to investigate a fundamental question: Is training PRMs necessary for effective code generation, or can outcomes alone provide sufficient supervision signals?
Experimental Design. Table 4 presents a comparison across two dimensions: granularity of supervision (line-level vs. outcome-level) and training approach (trained vs. inference-only). For line-level methods, the model generates step-by-step reasoning with numerical rewards assigned to each step, following mathematical reasoning approaches. For outcome-level methods, outcome rewards are computed after complete code generation and execution.
To determine whether the popular method of training line-level PRMs with higher-quality data can more effectively bridge the performance gap with our inference-only approach, we trained two PRMs with different data quality: PRM-GPT using a larger GPT-4o labeled dataset (containing 13644 synthetic labels for each line of reasoning text), and PRM-Human using human-filtered data where three authors spent 12 hours each to validate reasoning steps, achieving an inter-annotator agreement of 0.44 (Cohen’s Kappa) and extracting 836 high-quality steps for training. This experiment thus directly tests if providing PRMs with optimal, high-quality training data enables them to match or surpass the performance of our method.
Results and Analysis. The results confirm several key insights. First, outcome-level reward signals prove more effective than line-level signals, as line-level feedback can only evaluate incomplete thought processes with limited information. Second, and more importantly, our inference-only approach substantially outperforms all trained PRMs given the same calls, even when PRMs are trained with high-quality human annotations.
When combined with the computational cost analysis (Table 5), these results indicate that while training data quality does impact PRM performance, trained PRMs still underperform our hybrid approach of verifiable execution rewards combined with LLM self-critique. This suggests that outcomes alone, when properly integrated with LLM reasoning capabilities, provide superior supervision signals compared to learned reward models. The effectiveness stems from grounding supervision in concrete, verifiable execution feedback rather than learned approximations.
4.5 Computational Cost Analysis
A critical concern raised by reviewers is whether ORPS’s improvements come at the cost of increased computational overhead. To address this, we conduct controlled experiments limiting the total number of LLM calls, which is the bottleneck of inference-only methods, ensuring fair comparison under identical computational budgets.
Complexity Analysis. We analyzed the number of LLM calls by examining the source code of baseline methods. ORPS requires $2\times N\times(K\times T+1)$ calls, where $N$ is candidate samples, $K$ is beam size, and $T$ is reasoning steps. Reflexion (Shinn et al., 2024) requires $2\times T$ calls, while LDB (Zhong et al., 2024) requires $1+P\times T\times(2+B\times N)$ calls, where $P$ is pass_at_k, $B$ is code blocks, and $N$ is max trials. For trained PRMs, we replace self-critic LLM calls with PRM calls, maintaining identical call counts.
Results. Table 5 presents results on LBPP with Qwen-7B under controlled computational budgets. ORPS consistently outperforms all baselines across different call limits, demonstrating superior computational efficiency. Notably, repair-focused methods like REx (Tang et al., 2024a) show diminishing returns as compute increases, while ORPS scales effectively. These results confirm that ORPS achieves better performance through more effective utilization of computational resources, not brute-force computation.
Table 5: Computational Cost Analysis. Pass@1 performance on LBPP with Qwen-7B controlling LLM calls.
Method 20 Calls 50 Calls 100 Calls
Reflexion (2024) 37.0 40.7 39.5
LDB (2024) 37.0 36.4 37.0
REx (2024a) 43.2 53.7 54.3
PRM-GPT 44.4 37.0 35.8
PRM-Human 40.7 38.3 42.0
ORPS (Ours) 48.4 55.6 64.2
@@ 4.6 Scaling Analysis
While the previous analysis controlled LLM call budgets, we now examine performance scaling given the same number of generated code solutions. We compare ORPS against Best-of-N (BoN) sampling by varying candidate solutions on LBPP. The results in Figure 4 reveal distinct scaling behaviors: ORPS demonstrates superior scaling potential, with performance improving rapidly as more candidates become available. This suggests that the structured reasoning in our framework more effectively identifies and refines promising solutions from the candidate pool. In contrast, BoN exhibits slower improvements, indicating that simple selection from larger pools provides diminishing returns without strategic guidance.
Figure 4: Performance vs. Inference Budget. The y-axis represents Pass@1 scores on LBPP. The x-axis represents the number of candidates generated during inference.
@@ 4.7 Case Studies
Through detailed case studies, we demonstrate how our framework enhances code generation by improving reasoning. As shown in Appendix F, the response generated by the traditional CoT method for the Minimum Greatest Common Divisor problem in LBPP demonstrates that while the model provides a detailed thought process during solution generation, the complexity of the task results in an imperfect code implementation. For instance, in CoT’s approach, the reliance on nested loops and pairwise GCD calculations introduces inefficiencies and fails to address scalability for larger datasets. Similarly, our method’s initial implementation demonstrates a lack of robustness in handling edge cases and unnecessary redundancies in subset formation.
@@
Figure 5: Performance by Problem Class. Top-20 problem classes in LBPP showing success rates and unsolved cases for ORPS  vs baseline.
@@ 5 Conclusion
We introduce Outcome-Refining Process Supervision (ORPS), a unified framework that fundamentally challenges the necessity of specially trained Process Reward Models in code generation. By combining execution-driven feedback with LLMs’ intrinsic reasoning capabilities, ORPS demonstrates that verifiable outcomes can effectively guide process supervision without costly reward model training. Our comprehensive analysis across different data qualities—from GPT-4 labels to human annotations—shows that even high-quality trained PRMs underperform our inference-only approach, suggesting a paradigm shift in how we approach process supervision for complex reasoning tasks.
ORPS reveals several key insights: (1) Strategic reasoning over local repair: Unlike methods that incrementally fix code blocks, ORPS operates at a higher abstraction level, enabling exploration of fundamentally different algorithmic approaches rather than being trapped in local optima. (2) Reasoning space over model scale: Smaller models with ORPS often outperform larger variants using conventional methods, highlighting that structured reasoning frameworks can be more effective than pure parameter scaling. (3) Computational efficiency: ORPS achieves superior performance through strategic resource utilization rather than brute-force computation, scaling effectively while baselines show diminishing returns.
These findings collectively suggest that LLMs’ intrinsic capabilities, when properly guided by verifiable signals, can eliminate the need for costly reward model training. Future work could explore extending this framework to other domains requiring rigorous reasoning and verification, opening new avenues for developing more efficient and practical reasoning frameworks in large language models.
Impact Statement
This work advances code generation capabilities in large language models through more efficient reasoning processes. While our primary focus is methodological—improving algorithmic problem-solving without costly reward models—we acknowledge broader implications common to code generation systems. Enhanced programming assistants could democratize software development but may also lower barriers for generating malicious code. The framework’s emphasis on code efficiency could reduce computational overhead in generated programs, though its environmental impact depends on deployment contexts.
These considerations reflect well-established societal tradeoffs in AI-powered code tools rather than novel risks introduced by our approach. We encourage responsible deployment with standard safeguards against misuse, consistent with ethical practices for generative AI systems.
@@ A.1 Search Algorithm Hyperparameters (ORPS)
  • Process Reward Weight ($\alpha$): 0.5. This metric determines the proportion of the Process Reward in the step reward.
  • Outcome Reward Weight ($\beta$): 0.5. This metric determines the proportion of the Outcome Reward in the step reward.
Table 6: Performance Metrics Description. Our evaluation framework uses both dynamic execution profiling and static code analysis metrics to comprehensively assess code quality and efficiency.
Category Metric Description
Dynamic Execution Profiling
Time Enabled Total CPU time spent executing the code, measured in nanoseconds. Lower values indicate more efficient execution and better algorithmic optimization.
Instruction Count Number of CPU instructions executed during runtime. Reflects computational efficiency, with lower counts suggesting more optimized code paths and better algorithm implementation.
Branch Misses Frequency of incorrect branch predictions during execution. Lower values indicate better code predictability and CPU pipeline efficiency, resulting in faster execution times.
Page Faults Number of times the program needs to access virtual memory. Fewer page faults suggest better memory management and more efficient memory access patterns.
Static Analysis
Code Length Total number of lines in the source code. Generally, shorter code length indicates more concise solutions while maintaining readability and functionality.
AST Node Count Number of nodes in the Abstract Syntax Tree. Measures structural complexity of the code, with fewer nodes suggesting simpler and more maintainable implementation.
Cyclomatic Complexity Quantifies the number of linearly independent paths through the code. Lower values indicate easier-to-maintain and test code, reducing potential bug sources.
Cognitive Complexity Measures how difficult the code is to understand, based on control flow structures and nesting. Lower scores suggest more readable and maintainable code that is easier to debug.
@@ A.4 Model Training Configuration
  • Training Framework: llamafactory (Zheng et al., 2024b)
@@ A.5 Hardware Environment
Appendix B Additional Results on CodeContests
During rebuttal, reviewers were interested in an additional competitive programming contest benchmark, CodeContests (Li et al., 2022b). Due to page limitation, we include our new results here in Table 7.
Table 7: Pass@1 performance on CodeContests with Qwen-7B controlling LLM calls at 100.
Method Pass@1 Tests % Valid % Time (ms)
Reflexion 8.48 14.53 27.27 52318
LDB 7.88 11.79 21.21 3350
REx 13.33 20.17 32.12 27791
PRM-GPT 4.94 8.64 17.28 3084
PRM-Human 9.88 15.04 23.46 1985
ORPS (Ours) 20.61 28.36 40.61 36824
Appendix C AI Usage in Code Development
@@
Appendix D Impact of Optimizing Different Metrics on Code Quality
In ORPS, we integrate multiple evaluation metrics to guide the model’s reasoning and code generation. These metrics include both static analysis (e.g., AST nodes, cyclomatic complexity) and dynamic execution profiling (e.g., execution speed, CPU instruction count, branch mispredictions). To better understand their individual contributions, we conduct an ablation study where each metric is optimized in isolation.
D.1 Experimental Setup
To isolate the effect of each metric, we conduct a series of ablation experiments. In each experiment, the model receives rewards only from a single metric, while all other evaluation criteria remain unchanged. Specifically, we consider the following setups:
  • +AST Nodes: Encourages structurally simpler code by minimizing the number of AST nodes.
  • +Cyclomatic Complexity: Penalizes excessive branching and loop structures to improve maintainability.
  • +Cognitive Complexity: Rewards code that is easier to understand based on nested structures and control flow.
  • +Execution Speed: Optimizes for faster execution while maintaining correctness.
  • +CPU Instructions: Minimizes the number of CPU instructions executed.
  • +Branch Mispredictions: Encourages predictability to improve processor efficiency.
  • +Page Faults: Reduces memory access overhead for better performance.
  • +All Metrics: Incorporates all the above metrics into a single optimization objective.
D.2 Results and Discussion
As shown in Figure 6, optimizing for a single metric significantly improves performance in that specific aspect, yet comes at a severe cost to other dimensions of code quality. This phenomenon suggests that the model falls into a form of local metric optimality, where it overfits to the given reward signal while neglecting other critical properties of high-quality code.
More specifically, we observe that optimizing for static analysis metrics (e.g., AST nodes, cyclomatic complexity) often leads to a sharp decline in dynamic execution metrics (e.g., execution speed, CPU instructions). For instance, minimizing cyclomatic complexity encourages structurally simpler code, yet it may suppress more efficient algorithmic choices that involve loops and conditionals. Conversely, optimizing for execution speed often results in obfuscated or redundant code, as the model prioritizes raw performance over maintainability.
These observations highlight an inherent challenge in code optimization: static structure and dynamic efficiency often conflict when optimized in isolation. This suggests that achieving well-balanced code quality requires a multi-objective optimization strategy rather than single-metric reinforcement. Traditional reward models struggle in such scenarios, as they often assume reward signals are aligned across different dimensions. However, our results indicate that code generation requires more nuanced supervision—one that dynamically balances trade-offs between readability, maintainability, and execution efficiency.
Furthermore, the steep performance drop in non-optimized metrics suggests that reward sparsity is a critical issue in single-metric training. Since the model receives no information about other quality dimensions, it fails to generalize improvements beyond the specific reward it observes. This contrasts with human programming intuition, where engineers naturally balance multiple objectives, such as runtime efficiency, readability, and modularity. Future work could explore techniques like adaptive reward scaling, where the model dynamically adjusts its focus based on real-time trade-offs rather than rigid metric-specific optimization.
Figure 6: Impact of single-metric optimization on code quality. Each value represents the difference from the baseline. Optimizing for a single metric significantly improves performance in that dimension but leads to severe degradation in others.
Appendix E Correlation of Reasoning Quality and Code Quality
Our method is based on a core hypothesis: higher-quality reasoning leads to higher-quality code. To validate this assumption, we conducted a simple motivation experiment.
E.1 Experimental Setup
We first prompted GPT-4o to generate two different reasoning chains for each problem in the LBPP dataset, explicitly ensuring a clear quality difference between them. To verify that the generated reasoning indeed exhibited significant quality differences, we used GPT-4o again to reassess and confirm their relative quality.
Next, we concatenated each reasoning chain with the original problem description and fed them separately into GPT-4o to generate code solutions. Finally, we evaluated the quality of the generated code across multiple metrics.
E.2 Results and Analysis
As shown in Figure 7, all metrics are normalized, where higher values indicate better performance. The red bars represent the quality of code generated from higher-quality reasoning, while the blue bars correspond to the lower-quality reasoning.
From the results, we observe that better reasoning generally leads to better overall code quality, particularly in dynamic execution metrics. This suggests that high-quality reasoning not only improves correctness but also enhances execution efficiency, possibly by guiding the model to generate more optimal algorithmic structures.
Interestingly, the complexity of the generated code does not increase significantly. One possible explanation is that more detailed and higher-quality reasoning naturally guides the model toward generating more sophisticated solutions that handle a wider range of scenarios, rather than simply making the code structurally more complex.
These findings reinforce the importance of reasoning supervision in code generation. A strong reasoning framework helps models not only arrive at correct solutions but also optimize execution efficiency without unnecessarily inflating complexity.
Figure 7: Impact of reasoning quality on code quality. Each metric is normalized, where higher values indicate better performance. The red bars represent the code quality when generated from higher-quality reasoning, while the blue bars correspond to lower-quality reasoning. The results suggest that better reasoning significantly improves execution-related metrics while having a limited effect on structural complexity.
Appendix F Example Model Outputs